How do I solve this problem I learned PEMDA 2 weeks ago. I'm in the 5th grade, but not sure what the brackets mean. 9+5[4-(2+3)]... I had 14(1) but can't come up with the right answer.
Thanks

Answer by Alan3354(69443)

not sure what the brackets mean. 9+5[4-(2+3)]
---------
9+5[4-(2+3)]
= 9+5[4- 5]
= 9+5[-1]
= 9 - 5
= 4
--------
The brackets are the same as parentheses, or braces { }